OK, so I'm new here and have no clue what I'm doing, but I'll shoot.
I am completing the square, solving for X, and checking with the two solutions.
eg.X^2 + 2X + 3 = 0 Figuring out the problem, I got X = +/- i sqrt2.
For the check I used my calculator to do the solution into the problem with the + sign. But now I have tried doing it with the negative sign and it doesn't work. Need help bad!!!

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

X^2 + 2X + 3 = 0
X^2 + 2X = -3
x^2 + 2x + 1 = -2
(x+1)^2 = -2
x + 1 = +sqrt(-2), - sqrt(-2)
x = -1 + isqrt(2), -1 - isqrt(2)

Answer by ikleyn(52776)

.

On completing the square see the lesson
    - HOW TO complete the square - Learning by examples
in this site.

On completing the square to solve a quadratic equation see the lesson
    - HOW TO solve quadratic equation by completing the square - Learning by examples
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic  "Quadratic equations".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.